OpenStudy (anonymous):

concept of probabilty Let P(Z)=.40, P(Y)=.30, and P(Z∪Y)=.58 Find each of the given probability a,,P(Z'∩Y') @alexwee123

OpenStudy (anonymous):

@payin Are Are Z' and Y' the respective complements?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@genius12

OpenStudy (anonymous):

@payin

OpenStudy (anonymous):

I was able to derive that formula from the following:\[\bf P(Z' ∪ Y')=P(Z')+P(Y')-P(Z'∩Y')\]\[\bf \implies P(Z'∩Y')=P(Z')+P(Y')- P(Z' ∪ Y')\]

OpenStudy (anonymous):

@payin Do you see what I did?

OpenStudy (anonymous):

@payin Just a second I made a mistake. Let me correct myself.

OpenStudy (anonymous):

yeah i did but im tryna figure out how u got those figures

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

The formula that I just typed in is correct. I just plugged in the wrong values for P(Z'∪Y'). Let me type up the correct solution.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

We know P(Z) = 0.4 then P(Z') = 1 - 0.4 = 0.6 Similarly P(Y) = 0.3 then P(Y') = 1 - 0.3 = 0.7 Now saying P(Z'∩Y') is the same as saying probability of Z' and Y' which is:\[\bf P(Z'∩Y')=0.6 \times 0.7 = ?\] @payin

OpenStudy (dan815):